# Curating art exhibitions using machine learning

### v1

@@ 3.4 Full OpenAI approach including embedding and mapping
Fourth and finally, we fully use external help, this time by using one of the latest OpenAI GPT (Chat Generative Pre-Trained Transformer models – Radford et al. (2018); Brown et al. (2020)), in this case gpt-4o-mini. These GPT-based models are trained from very large amounts of data scraped from the open Internet, in particular Wikipedia, the arXiv.org e-Print archive, Reddit, among others. These models are a subset of the so-called large language models (LLMs), whereby one maps an input variable $x$, corresponding to the question, to an output variable $y$, corresponding to the answer, using a complex mathematical function, which is encoded in the values of the weights of the nodes of a (very large [^14] It is estimated that this particular model may have up to 8 billion parameters. However, this is just an estimate as OpenAI usually does not officially disclose exact number of parameters in their models.) neural network. The transformation of the question, which is a plain text input, into a numerical vector $x$ of very high dimension is similar to the method described in the Sections 3.1, 3.2 and 3.3, an embedding, but it is assumed (since again it is not disclosed by OpenAI) that gpt-4o-mini has its own embedding model. The same goes for the output variable $y$. The mapping is not an explicit analytical equation, but rather a numerical function discovered by the training process using a technique called back-propagation (Rumelhart et al., 1986), and encoded in the numerical weights of the nodes of a very large neural network. In order to use gpt-4o-mini to create a list of artworks for an exhibition, we use the following steps. First, we fine-tune the model, by showing it examples in the form $\langle x,y\rangle$, where $x$ is the input text as in the listing above 2 and $y$ is a table of the metadata items for all the artworks presented within the exhibition (see Listing 4 for an example). Fine-tuning is a machine learning technique whereby one takes an already pre-trained model (such as one of the GPT models) and adapts it to a new domain and/or makes it more specific to a sub-domain (Yosinski et al., 2014; Howard and Ruder, 2018; Radford et al., 2018; Wei et al., 2021). The fine-tuning process re-trains the last few layers of the neural network, while maintaining most layers intact. This process then improves the model for the new domain and/or sub-domain we are interested in. [^15] It is well known that there is a kind of feature hierarchy in neural networks, whereby the first layers encode small features while the last layers encode overall generic task or domain specific features, as demonstrated in Yosinski et al. (2014). A common analogy used is of a very young child learning a second language, the knowledge gained during the first language acquisition is re-used and fine-tuned for the second language. An example would be to retrain the weights between the second hidden layer and the output layer on our example toy model neural network in Figure 1.
@@
[…]
Figure 19: Two steps neural network approach using the fine-tune versions of OpenAI’s gpt-4o-mini (version 2024-07-18). The input text is the title/description of each exhibition, and the output is a formatted version of the listing 4.
After we fine-tune the model, in this case gpt-4o-mini (version 2024-07-18), we can use it directly. We can query using inputs within the validation set or out-of-sample test set examples. As before, we split the set of exhibitions we have into a training and a validation set, in the proportion $80\%$ examples for training (a total of 188 examples) and $20\%$ samples for validation or in-sample testing (a total of 48 examples). We notice that the output of the model this time can be of variable length, that is to say, the number of predicted artworks this time is variable (in all the other models before, the selection is done based on the number of actual artworks in each exhibition). This we see as an advantage of this model. We also notice that more often than not the model outputs a series of artworks with generalised tags in an invalid format (e.g., values missing, rows/columns incomplete, etc.). In these, we simply rerun the model until we get a output that is fully capable of being parsed, having given up trying to fix the invalid formats. We also notice that the output of GPT although it is a list of rows with “Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”, these are not necessary either real existing artworks (i.e., inconsistent data can be output, the so-called LLM hallucinations; see e.g., Farquhar et al. (2024) and references therein) or artworks within the Met Museum. Therefore, we still need to use the Met Museum database to find the nearest artworks to our output, using the same approach with the hit function as defined in Equation 1.
We plot below the MSE of the training and validation sets for our fine-tuning in Figure 20. This time the convergence was not as clear as in the other models, however, this was not a problem as we can see later – fine-tuning typically changes the model weights only slightly rather than completely refitting the training dataset, so loss (MSE) reduction can be modest, see e.g., Howard and Ruder (2018, Fig. 4, p. 336).
@@
Figure 21: Marble composite of images from the Met Museum corresponding to the run for the fine-tune model trained on the training set only, using as input the text (title/description) of the exhibition and as output the actual list of artworks and their metadata in their original format as in the listing 4. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API (The Metropolitan Museum of Art, 2025a). Non-available images are marked by their object id JPEG name.
@@ 4 Possible further research approaches
Second, we could also enhance the number of usable exhibitions by adding those that are composed of artworks that are lent to the Met Museum. However, the reason those were removed is that the Met Museum database does not contain those artworks or theirs associated metadata. To manually add these would be a possible but very difficult and time-consuming task. Third, we could also gather data from other museums, and merge with the dataset we already have from the Met Museum. There are many possibilities, e.g., the Museum of Modern Art (The Museum of Modern Art, 2025a) provides a collection API; the British Museum has an extensive collection of more than five million objects (British Museum, 2025); the National Gallery in London also provides an API (The National Gallery, 2025); and many others. However, this would be a much larger task, and one would have to use as input not just the exhibition test text but also the target museum, as one would need constraints to minimise artworks hosted outside this target museum, to make the exhibition as feasible as possible. Furthermore, each museum uses its own format for the artwork metadata, with different field names/meanings, making the merging of data across multiple sources a very difficult and subjective task.
@@
In terms of modelling itself, we can also think of four different approaches we could use. First, as mentioned in Section 3.1 we could use a Long Short-Term Memory (LSTM) to turn our task into a variable input/variable output problem. This is similar to what is used in human languages translations, e.g., English text to a French text (Sutskever et al., 2014). The advantage of a LSTM neural network is that being a recurrent neural network it would allow for variable input/output sizes, which is part of our problem. While we have gone around this issue in this article by constructing models that transform the problem into a fixed input/output size, there is in principle no actual obstacle in using a LSTM to output a variable list of artworks. However, in practice with such a small amount of training examples (maximum of 236), the data size made this approach unrealistic. Maybe with a much larger dataset, including other museum’s data would make it possible. Second, we could construct a slightly different approach as we used in model in Section 3.3. While in that model we created an embedding of the input exhibition title/description and an embedding of the output concatenated list of metadata generalised tags (one string), we could instead unfold our training dataset by mapping each exhibition’s embedding to each metadata string embedding, so e.g., if one exhibition had $n$ artworks each with $m$ generalised tags, then we would create in effect $nm$ input/output embeddings. The reason we did not try this is that the resulting graph database and subsequent search would be too large/slow for our purpose, given our computational constrains. Third, we could in principle add our own tags by downloading all images of each artwork from the Met Museum database and using a Convolutional Neural Network (CNN) to extract our own extra features/metadata (LeCun et al., 2002; Krizhevsky et al., 2012). However, this is a massive undertaking, requiring lots of computational power and storage space. Furthermore, this would require inside access as most images are not in the public domain. Of the total $484,956$ works of art, only $248,472$ are in the public domain and available for download. Fourth, and finally, we could also include in our models more of the Met Museum generalised tags. There are 50+ fields that the Met Museum provides in its full artwork metadata set, we just use 6 of those, as we recall these were “Department”,“Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”. We did attempt to add other fields, but saw such minor or null improvements, at a computational cost and reduction of simplicity/understanding which we deemed unnecessary.
@@ 5 Conclusion
We have attempted to create four machine learning or artificial intelligence models that learn the behaviour of human art curators. We used data from 25 years of Metropolitan Museum of Art in New York to train and optimise the models. Three of the four models can be shown to imitate human behaviour to a reasonable accuracy, namely, well above random choice for the set of existing exhibitions and the set of existing artworks in the full collection. Furthermore, as the models get increasingly complex, i.e., as we move from relying simply on the dataset statistical information and start using more and more the knowledge inherent in the so-called large language embedding and query models, we can show we get some improvement in the accuracy, measured by the selection of artworks and its metadata. Nonetheless, the improvement is not very large, and this seems to show that even models with modest size (e.g., our second model with input embeddings and output metadata probabilities) can perform almost as well as very large models using brute force. In particular, this can be seen in our out-of-sample experiment (the one with the impressionists related input text). Finally, we believe that as we gather more data in the future and improve the models, that these kinds of machine learning approaches may be able to get closer and closer to the level of curatorship that humans possess, although it remains a subject of discussion in philosophy of art if there is any artistic value in this. This could be of interest to art galleries who wish to use AI models to autonomously curate their in-house exhibitions and/or to allow the public to access a website or an app that would create virtual exhibitions based on the public’s input exhibition title and description.
@@
  • Conflict of interest or competing interests: None;

### v2

@@ 3.4 Full OpenAI approach including embedding and mapping
Fourth and finally, we fully use external help, this time by using one of the latest OpenAI GPT (Chat Generative Pre-Trained Transformer models – Radford et al. [2018], Brown et al. [2020]), in this case gpt-4o-mini. These GPT-based models are trained from very large amounts of data scraped from the open Internet, in particular Wikipedia, the arXiv.org e-Print archive, Reddit, among others. These models are a subset of the so-called large language models (LLMs), whereby one maps an input variable $x$, corresponding to the question, to an output variable $y$, corresponding to the answer, using a complex mathematical function, which is encoded in the values of the weights of the nodes of a (very large [^14] It is estimated that this particular model may have up to 8 billion parameters. However, this is just an estimate as OpenAI usually does not officially disclose exact number of parameters in their models.) neural network. The transformation of the question, which is a plain text input, into a numerical vector $x$ of very high dimension is similar to the method described in the Sections 3.1, 3.2 and 3.3, an embedding, but it is assumed (since again it is not disclosed by OpenAI) that gpt-4o-mini has its own embedding model. The same goes for the output variable $y$. The mapping is not an explicit analytical equation, but rather a numerical function discovered by the training process using a technique called back-propagation [Rumelhart et al., 1986], and encoded in the numerical weights of the nodes of a very large neural network. In order to use gpt-4o-mini to create a list of artworks for an exhibition, we use the following steps. First, we fine-tune the model, by showing it examples in the form $\langle x,y\rangle$, where $x$ is the input text as in the listing above 2 and $y$ is a table of the metadata items for all the artworks presented within the exhibition (see Listing 4 for an example). Fine-tuning is a machine learning technique whereby one takes an already pre-trained model (such as one of the GPT models) and adapts it to a new domain and/or makes it more specific to a sub-domain [Yosinski et al., 2014, Howard and Ruder, 2018, Radford et al., 2018, Wei et al., 2021]. The fine-tuning process re-trains the last few layers of the neural network, while maintaining most layers intact. This process then improves the model for the new domain and/or sub-domain we are interested in. [^15] It is well known that there is a kind of feature hierarchy in neural networks, whereby the first layers encode small features while the last layers encode overall generic task or domain specific features, as demonstrated in Yosinski et al. [2014]. A common analogy used is of a very young child learning a second language, the knowledge gained during the first language acquisition is re-used and fine-tuned for the second language. An example would be to retrain the weights between the second hidden layer and the output layer on our example toy model neural network in Figure 1.
@@
Titles andoverview texts+artworks and metadataFine-tuningOpenAIgpt-4o-miniFine-tunedmodelTitle andoverview textFine-tunedmodelPredictedArtworks andmetadata
Figure 19: Two-step neural network approach using the fine-tune versions of OpenAI’s gpt-4o-mini (version 2024-07-18). The input text is the title/description of each exhibition, and the output is a formatted version of the listing 4.
After we fine-tune the model, in this case gpt-4o-mini (version 2024-07-18), we can use it directly. We can query using inputs within the validation set or out-of-sample test set examples. As before, we split the set of exhibitions we have into a training and a validation set, in the proportion $80\%$ examples for training (a total of 188 examples) and $20\%$ samples for validation or in-sample testing (a total of 48 examples). We notice that the output of the model this time can be of variable length, that is to say, the number of predicted artworks this time is variable (in all the other models before, the selection is done based on the number of actual artworks in each exhibition). This we see as an advantage of this model. We also notice that more often than not the model outputs a series of artworks with generalised tags in an invalid format (e.g., values missing, rows/columns incomplete, etc.). In these, we simply rerun the model until we get a output that is fully capable of being parsed, having given up trying to fix the invalid formats. We also notice that the output of GPT although it is a list of rows with “Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”, these are not necessary either real existing artworks (i.e., inconsistent data can be output, the so-called LLM hallucinations; see e.g., Farquhar et al. [2024] and references therein) or artworks within the Met Museum. Therefore, we still need to use the Met Museum database to find the nearest artworks to our output, using the same approach with the hit function as defined in Equation 1.
We plot the MSE of the training and validation sets for our fine-tuning in Figure 20. This time the convergence was not as clear as in the other models, however, this was not a problem as we can see later – fine-tuning typically changes the model weights only slightly rather than completely refitting the training dataset, so loss (MSE) reduction can be modest, see e.g., Howard and Ruder [2018, Fig. 4, p. 336].
@@
Figure 21: Marble composite of images from the Met Museum corresponding to the run for the fine-tune model trained on the training set only, using as input the text (title/description) of the exhibition and as output the actual list of artworks and their metadata in their original format as in the listing 4. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API [The Metropolitan Museum of Art, 2025a]. Non-available images are marked by their object id JPEG name.
@@ 4 Possible further research approaches
Second, we could also enhance the number of usable exhibitions by adding those that are composed of artworks that are lent to the Met Museum. However, the reason those were removed is that the Met Museum database does not contain those artworks or theirs associated metadata. To manually add these would be a possible but very difficult and time-consuming task. Third, we could also gather data from other museums, and merge with the dataset we already have from the Met Museum. There are many possibilities, e.g., the Museum of Modern Art [The Museum of Modern Art, 2025a] provides a collection API; the British Museum has an extensive collection of more than five million objects [British Museum, 2025]; the National Gallery in London also provides an API [The National Gallery, 2025]; and many others. However, this would be a much larger task, and one would have to use as input not just the exhibition test text but also the target museum, as one would need constraints to minimise artworks hosted outside this target museum, to make the exhibition as feasible as possible. Furthermore, each museum uses its own format for the artwork metadata, with different field names/meanings, making the merging of data across multiple sources a very difficult and subjective task.
@@
In terms of modelling itself, we can also think of four different approaches we could use. First, as mentioned in Section 3.1 we could use a Long Short-Term Memory (LSTM) to turn our task into a variable input/variable output problem. This is similar to what is used in human languages translations, e.g., English text to a French text [Sutskever et al., 2014]. The advantage of a LSTM neural network is that being a recurrent neural network it would allow for variable input/output sizes, which is part of our problem. While we have gone around this issue in this article by constructing models that transform the problem into a fixed input/output size, there is in principle no actual obstacle in using a LSTM to output a variable list of artworks. However, in practice with such a small amount of training examples (maximum of 236), the data size made this approach unrealistic. Maybe with a much larger dataset, including other museum’s data would make it possible. Second, we could construct a slightly different approach as we used in model in Section 3.3. While in that model we created an embedding of the input exhibition title/description and an embedding of the output concatenated list of metadata generalised tags (one string), we could instead unfold our training dataset by mapping each exhibition’s embedding to each metadata string embedding, so e.g., if one exhibition had $n$ artworks each with $m$ generalised tags, then we would create in effect $nm$ input/output embeddings. The reason we did not try this is that the resulting graph database and subsequent search would be too large/slow for our purpose, given our computational constrains. Third, we could in principle add our own tags by downloading all images of each artwork from the Met Museum database and using a Convolutional Neural Network (CNN) to extract our own extra features/metadata [LeCun et al., 2002, Krizhevsky et al., 2012]. However, this is a massive undertaking, requiring lots of computational power and storage space. Furthermore, this would require inside access as most images are not in the public domain. Of the total $484,956$ works of art, only $248,472$ are in the public domain and available for download. Fourth, and finally, we could also include in our models more of the Met Museum generalised tags. There are 50+ fields that the Met Museum provides in its full artwork metadata set, we just use 6 of those, as we recall these were “Department”,“Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”. We did attempt to add other fields, but saw such minor or null improvements, at a computational cost and reduction of simplicity/understanding which we deemed unnecessary.
@@ 5 Conclusion
We have attempted to create four machine learning or artificial intelligence models that learn the behaviour of human art curators. We used data from 25 years of exhibitions at the Metropolitan Museum of Art in New York to train and optimise the models. Three of the four models can be shown to imitate human behaviour to a reasonable accuracy, namely, well above random choice for the set of existing exhibitions and the set of existing artworks in the full collection. Furthermore, as the models get increasingly complex, i.e., as we move from relying simply on the dataset statistical information and start using more and more the knowledge inherent in the so-called large language embedding and query models, we can show we get some improvement in the accuracy, measured by the selection of artworks and its metadata. Nonetheless, the improvement is not very large, and this seems to show that even models with modest size (e.g., our second model with input embeddings and output metadata probabilities) can perform almost as well as very large models using brute force. In particular, this can be seen in our out-of-sample experiment (the one with the impressionists related input text). Finally, we believe that as we gather more data in the future and improve the models, that these kinds of machine learning approaches may be able to get closer and closer to the level of curatorship that humans possess, although it remains a subject of discussion in philosophy of art if there is any artistic value in this. This could be of interest to art galleries who wish to use AI models to autonomously curate their in-house exhibitions and/or to allow the public to access a website or an app that would create virtual exhibitions based on the public’s input exhibition title and description.
@@
  • The author declares no existing conflicts of interest;